Use the explicit formula an=a1+(n-1)d to find the 1000th term of the sequence below.
24,31,38,45,52,..

Answer by stanbon(75887)

Use the explicit formula an=a1+(n-1)d to find the 1000th term of the sequence below.
24,31,38,45,52,..
00000
a(1000) = 24 + 999*7 = 7017
================
Cheers,
Stan H.
================